1 kg of ait at 1 bar 15 degree centridage is compressed reversibly and adiabtically in a pressure of 4 bar.Calculate the final temperature and the work input.


1
Expert's answer
2021-04-12T01:53:27-0400

Adiabatic Compression

"\gamma" for adiabatic air process = 1.4

T2/T1 = (P2/P1)(1.4 - 1)/1.4

T2 = (15+273 K)(4 x 105 Pa / 1 x 105 Pa)0.4/1.4 = 428 K

W = (P1V1 - P2V2) / ("\gamma" - 1) = R(T2 - T1) / ("\gamma" - 1)

W = 287 J/kg K (288 - 428 K) x 1 kg / (1.4 - 1) = -100.45 kJ